1) Hydrogen iodide 0.2 moles each of H2 and I2 into a 1litre flask. Calculate the equilibrium contact value (Kc) for the reaction when a) 0.0z
7 moles per litre eaach of H2 and I2 was present at equilibrium. B) 0.20 moles per litre each of H2 and I2 had each at the time equilibrium was established.
1
Expert's answer
2013-05-27T12:58:04-0400
The reaction is:
H2 + I2 = 2HI

A)
The concentrations are:
C(H2) = 7 / 1 = 7 M
C(I2) = 7/ 1 = 7 M
C(HI) = 7*2 / 1 = 14 M
Kc = [HI]2/[H2][I2]
so the equilibrium contact value will be
Kc = 142 / 7*7 = 4

B)
The concentrations are:
C(H2) = 0.2 / 1 = 0.2 M
C(I2) = 0.2 / 1 = 0.2 M
C(HI) = 0.2 *2 / 1 = 0.4 M
Kc = [HI]2/[H2][I2]
so the equilibrium contact value will be
Kc = 0.42 / 0.2*0.2 = 4
Answer: 4